### v1

# Gaussian periods and Shanks’ cubic polynomials. II

Abstract.
We give a linear relation between a cubic Gaussian period and a root of Shanks’ cubic polynomial in wildly ramified cases.
Key words and phrases:
Gaussian period, period polynomial, Shanks’ cubic polynomial
2020 Mathematics Subject Classification:
Primary 11R04, 11R16, Secondary 11C08, 11R80.
This work was supported by JSPS KAKENHI Grant Number JP21K03181
1. Introduction
Let $L$ be a cyclic cubic field. The conductor $\mathfrak{f}$ of $L$ should be
(1.1) $\mathfrak{f}=\begin{cases}p_{1}\cdots p_{\nu}&\text{if $3\nmid\mathfrak{f}$\, (tamely ramified)},\\ 3^{2}p_{1}\cdots p_{\nu}&\text{if $3|\mathfrak{f}$\, (wildly ramified)},\end{cases}$
where $p_{1},\ldots,p_{\nu}$ are different prime numbers satisfying $p_{1}\equiv\cdots\equiv p_{\nu}\equiv 1\pmod{3}$ ([9, p. 10]). For a positive integer $n$, let $\zeta_{n}=e^{2\pi i/n}$ be the $n$-th root of unity. We define the Gaussian periods $\eta_{i}\,(i=0,1,2)$ of $L$ by
(1.2) $\eta_{0}={\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L}(\zeta_{\mathfrak{f}}),\quad\eta_{1}=\sigma(\eta_{0}),\quad\eta_{2}=\sigma^{2}(\eta_{0})$
where ${\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L}$ is the trace map from $\mathbb{Q}(\zeta_{\mathfrak{f}})$ to $L$, and $\sigma$ is a generator of ${\rm Gal}(L/\mathbb{Q})$. It is known that $L=\mathbb{Q}(\eta_{i})$ holds for any $i=0,1,2$. We define the period polynomial $P(X)\,(\in\mathbb{Z}[X])$ by
(1.3) $P(X)=(X-\eta_{0})(X-\eta_{1})(X-\eta_{2}).$
An explicit formula of $P(X)$ is known as follows ([6], [9], [7, p. 90], [14, p. 8–9]).
[…]
where $\mu$ is the Möbius function, and $M\,(\in\mathbb{Z})$ satisfies the following for some $N\,(\in\mathbb{Z})$.
(1.5) $4\mathfrak{f}=M^{2}+27N^{2},$
[…]
On the other hand, for an integer $\mathfrak{f}$ given in the form (1.1), there are exactly $2^{\nu-1}$ (resp. $2^{\nu}$) pairs $(M,N)\in\mathbb{Z}\times\mathbb{Z}$ which satisfy (1.5) ([5, p. 342-343,  p. 364  Exercise 18]), and each pair $(M,N)$ corresponds to exactly $2^{\nu-1}$ (resp. $2^{\nu}$) cyclic cubic fields $\mathbb{Q}(\eta_{0})$ with conductor $\mathfrak{f}$ in the case of tamely (resp. wildly) ramified.
Next, we will explain the known results on the connection between the period polynomial and Shanks’ cubic polynomial. For $n\in\mathbb{Q}$, we define Shanks’ cubic polynomial $f_{n}(X)\,(\in\mathbb{Q}[X])$ by
(1.6) $f_{n}(X)=X^{3}-nX^{2}-(n+3)X-1.$
The discriminant of $f_{n}(X)$ is $d(f_{n})=(n^{2}+3n+9)^{2}$. For a root of $\rho_{n}$ of $f_{n}(X)$, let $L_{n}=\mathbb{Q}(\rho_{n})$ and $G={\rm Gal}(L_{n}/\mathbb{Q})$. If $f_{n}(X)$ is irreducible over $\mathbb{Q}$, then $L_{n}$ is a cyclic cubic field. In this case, we put $\rho_{n}^{\prime}=\sigma(\rho_{n})$ and $\rho_{n}^{\prime\prime}=\sigma^{2}(\rho_{n})$ where $\sigma$ is a generator of $G$. It is known that $f_{n}(X)$ is a generic cyclic cubic polynomial ([17, chap. 1]). Namely, for any cyclic cubic field $L$, there exists $n\in\mathbb{Q}$ such that $L=L_{n}$. If $n\in\mathbb{Z}$, then $f_{n}(X)$ is always irreducible, and the field $L_{n}$ is known as the simplest cubic field (see [18], [19]).
As explained below, in special cases, the relation between Shanks’ polynomial $f_{n}(X)$ and the period polynomial $P(X)$ is known. Let $\mathfrak{f}=p_{1}\cdots p_{\nu}$ be an integer where $p_{1},\ldots,p_{\nu}$ are different prime numbers which satisfy $p_{1}\equiv\cdots\equiv p_{\nu}\equiv 1\pmod{3}$, and a pair $(M,N)\in\mathbb{Z}\times\mathbb{Z}$ satisfies (1.5). Assume that $N=1$. In this case, it is known that $L=\mathbb{Q}(\eta_{0})$ is a simplest cubic field $L_{n}$ for $n=\left(M-3\right)/2\in\mathbb{Z}$, and there is a linear relation between the Gaussian period $\eta_{i}$ and a root of $f_{n}(X)$ ([11, p. 536], [4], [10, Proposition 2.2]). We can easily check $4\mathfrak{f}=M^{2}+27=4(n^{2}+3n+9)$, and hence $\mathfrak{f}=n^{2}+3n+9$. If we use the explicit formula (1.4) of the period polynomial, we can check
(1.7) $\mu(\mathfrak{f})f_{n}(X)=P\left(\mu(\mathfrak{f})(X+v_{n})\right)$
where $v_{n}=\left(1-n\right)/3\,(\in\mathbb{Z})$, and hence we obtain
(1.8) $\{\eta_{0},\eta_{1},\eta_{2}\}=\left\{\mu(\mathfrak{f})(\rho_{n}+v_{n}),\mu(\mathfrak{f})(\rho_{n}^{\prime}+v_{n}),\mu(\mathfrak{f})(\rho_{n}^{\prime\prime}+v_{n})\right\}$
and $L=L_{n}$.
In this paper, we will extend these results (1.7) and (1.8) for general pairs $(M,N)$ (not necessarily $N=1$) of (1.5) without the explicit formula (1.4) of $P(X)$ in the case of wildly ramified. In the case of tamely ramified, the author gave the following theorem.
Theorem 1 ([3]).
Let $\mathfrak{f}=p_{1}\cdots p_{\nu}$ be an integer where $p_{1},\ldots,p_{\nu}$ are different prime numbers which satisfy $p_{1}\equiv\cdots\equiv p_{\nu}\equiv 1\pmod{3}$, and a pair $(M,N)\in\mathbb{Z}\times\mathbb{Z}$ satisfies (1.5). Put $n_{1}=(M-3N)/2,\ n_{2}=N$ and $n=n_{1}/n_{2}$. Then $n_{1}$ and $n_{2}$ satisfy the following.
  1. (1)
    $n_{1}$ and $n_{2}$ are coprime integers.
  2. (2)
    $n_{1}^{2}+3n_{1}n_{2}+9n_{2}^{2}=\mathfrak{f}$.
  3. (3)
    $f_{n}(X)$ is irreducible and the conductor of the cyclic cubic field of $L_{n}=\mathbb{Q}(\rho_{n})$ is $\mathfrak{f}$.
  4. (4)
    $n_{2}^{3}\mu(\mathfrak{f})f_{n}(X)=P\left(\mu(\mathfrak{f})\left(n_{2}X+\frac{1-n_{1}}{3}\right)\right)$ holds, where $P(X)$ is the period polynomial given by (1.3) whose roots are the Gaussian periods $\eta_{0},\eta_{1},\eta_{2}$ of $L_{n}$.
  5. (5)
    $\{\eta_{0},\eta_{1},\eta_{2}\}=\left\{\mu(\mathfrak{f})\left(n_{2}\rho_{n}+\frac{1-n_{1}}{3}\right),\mu(\mathfrak{f})\left(n_{2}\rho^{\prime}_{n}+\frac{1-n_{1}}{3}\right),\mu(\mathfrak{f})\left(n_{2}\rho^{\prime\prime}_{n}+\frac{1-n_{1}}{3}\right)\right\}$.
Furthermore, all cyclic cubic fields with conductor $\mathfrak{f}$ are given by $L_{n}$ for such $n=n_{1}/n_{2}$.
We will give a similar theorem in the case of wildly ramified. The main result of this paper is as follows.
Theorem 2.
Let $\mathfrak{f}=3^{2}p_{1}\cdots p_{\nu}$ be an integer where $p_{1},\ldots,p_{\nu}$ are different prime numbers which satisfy $p_{1}\equiv\cdots\equiv p_{\nu}\equiv 1\pmod{3}$, and a pair $(M,N)\in\mathbb{Z}\times\mathbb{Z}$ satisfies (1.5). Put $n_{1}=(M-3N)/2,\ n_{2}=N$ and $n=n_{1}/n_{2}$. Then $n_{1}$ and $n_{2}$ satisfy the following.
  1. (1)
    $n_{1}$ and $n_{2}$ are coprime integers.
  2. (2)
    $n_{1}^{2}+3n_{1}n_{2}+9n_{2}^{2}=\mathfrak{f}$.
  3. (3)
    $f_{n}(X)$ is irreducible and the conductor of the cyclic cubic field of $L_{n}=\mathbb{Q}(\rho_{n})$ is $\mathfrak{f}$.
  4. (4)
    $n_{2}^{3}\mu(\mathfrak{f}/9)f_{n}(X)=P\left(\mu(\mathfrak{f}/9)\left(n_{2}X-\frac{n_{1}}{3}\right)\right)$ holds, where $P(X)$ is the period polynomial given by (1.3) whose roots are the Gaussian periods $\eta_{0},\eta_{1},\eta_{2}$ of $L_{n}$.
  5. (5)
    $\{\eta_{0},\eta_{1},\eta_{2}\}=\left\{\mu(\mathfrak{f}/9)\left(n_{2}\rho_{n}-\frac{n_{1}}{3}\right),\mu(\mathfrak{f}/9)\left(n_{2}\rho^{\prime}_{n}-\frac{n_{1}}{3}\right),\mu(\mathfrak{f}/9)\left(n_{2}\rho^{\prime\prime}_{n}-\frac{n_{1}}{3}\right)\right\}$.
Furthermore, all cyclic cubic fields with conductor $\mathfrak{f}$ are given by $L_{n}$ for such $n=n_{1}/n_{2}$.
We will prove the theorem without known result (1.4) on the period polynomial, and use recent results [2] on the Galois module structure of the ring of integers of cyclic cubic fields.
Remark 1.
  1. (1)
    If we use the explicit formula (1.4) of $P(X)$, then we can easily prove (4) and (5) by direct calculation. Conversely, if we have (4) and (5), then we can obtain the explicit formula (1.4).
  2. (2)
    By the theorem, we know that if $L$ is a wildly ramified cubic field, then there exists coprime integers $n_{1}$ and $n_{2}$ which satisfy $L=L_{n}$ for $n=n_{1}/n_{2}$ and $(n_{1}^{2}+3n_{1}n_{2}+9n_{2}^{2})/9$ is square-free.
  3. (3)
    Let $M$ and $N$ be integers satisfying (1.5) and put
    (1.9) $\dfrac{M+3N\sqrt{-3}}{2}=\begin{cases}\pi_{1}\cdots\pi_{\nu}&\text{if $3\nmid\mathfrak{f}$},\\ 3\zeta_{3}^{\pm 1}\pi_{1}\cdots\pi_{\nu}&\text{if $3|\mathfrak{f}$},\end{cases}$
    where $\pi_{i}\in\mathbb{Z}[\zeta_{3}]$ are prime elements which divide $p_{i}$ and $\tau(\chi_{p_{i}})^{3}=p_{i}\pi_{i}$ for the character $\chi_{p_{i}}$ defined by $\chi_{p_{i}}(a)\equiv a^{(p_{i}-1)/3}\pmod{(\pi_{i})}$, and ${\tau(\chi_{p_{i}})=\sum_{a\in(\mathbb{Z}/p_{i}\mathbb{Z})^{\times}}\chi_{p_{i}}(a)\zeta_{p_{i}}^{a}}$ is the Gaussian sum. Let $\chi_{3^{2}}$ be the character defined by $\chi_{3^{2}}(a)=\zeta_{3}^{\pm(a^{2}-1)/3}$ (double sign in same order in (1.9)). Put
    $\chi=\begin{cases}\chi_{p_{1}}\cdots\chi_{p_{\nu}}&\text{if $3\nmid\mathfrak{f}$},\\ \chi_{3^{2}}\chi_{p_{1}}\cdots\chi_{p_{\nu}}&\text{if $3|\mathfrak{f}$}.\end{cases}$
    Then the cyclic cubic fields $L_{n}$ of Theorems 1 and 2 are the fields corresponding to ${\rm Ker}\,\chi\leq(\mathbb{Z}/\mathfrak{f}\mathbb{Z})^{\times}$ ([9, p. 12–13]).
2. Preliminaries
In this section, we prove some lemmas and a theorem used in the proof of Theorem 2.
Lemma 1.
Let $n=n_{1}/n_{2}$ be a rational number where the integers $n_{1}$ and $n_{2}$ are coprime. Suppose that $3|n_{1},9||\Delta_{n}$ and $\Delta_{n}/9$ is square-free, where $\Delta_{n}=n_{1}^{2}+3n_{1}n_{2}+9n_{2}^{2}$. Then the cubic polynomial $f_{n}(X)$ is irreducible over $\mathbb{Q}$.
Proof.
First, we show that $\Delta_{n}/9$ and $2n_{2}+3n_{2}$ are coprime. Let $p$ be a prime number which divides both $\Delta_{n}/9$ and $2n_{1}+3n_{2}$. We can easily check $p\neq 2,3$ since $2\nmid(\Delta_{n}/9)$ and $3\nmid(\Delta_{n}/9)$. Furthermore, since $4\Delta_{n}=(2n_{1}+3n_{2})^{2}+27n_{2}^{2}$ and $p\neq 3$, we have $p|n_{2}$. Hence we have $p|2n_{1}$ since $p$ divides both $2n_{1}+3n_{2}$ and $n_{2}$. This is a contradiction since $p\neq 2$ and $(n_{1},n_{2})=1$. Therefore $\Delta_{n}/9$ and $2n_{1}+3n_{2}$ are coprime. The irreducibility of $f_{n}(X)$ can be obtained by using Eisenstein’s criterion for the right-hand side of
$(3n_{2})^{3}f_{n}\left(\frac{X}{3n_{2}}+\frac{n}{3}\right)=X^{3}-3\Delta_{n}X-(2n_{1}+3n_{2})\Delta_{n}$
and a prime factor of $\Delta_{n}/9$ if $\Delta_{n}/9\neq 1$. If $\Delta_{n}/9=1$, then we have $n_{2}=\pm 1$ since $1=\Delta_{n}/9=(n_{1}/3+n_{2}/2)^{2}+3/4\,n_{2}^{2}\geq 3/4\,n_{2}^{2}$, and hence $n=n_{1}/n_{2}\in\mathbb{Z}$ and $f_{n}(x)$ is irreducible over $\mathbb{Q}$ (in this case, we have n=0,3n=0,-3 and $L_{0}=L_{-3}$). ∎
Lemma 2.
Let $n=n_{1}/n_{2}$ and $n^{\prime}=n_{1}^{\prime}/n_{2}^{\prime}$ be rational numbers where the integers $n_{1},n_{2}$ and $n_{1}^{\prime},n_{2}^{\prime}$ satisfy $(n_{1},n_{2})=(n_{1}^{\prime},n_{2}^{\prime})=1$. Put $\Delta_{n}=n_{1}^{2}+3n_{1}n_{2}+9n_{2}^{2}$ and $\Delta_{n^{\prime}}={n_{1}^{\prime}}^{2}+3n_{1}^{\prime}n_{2}^{\prime}+9{n_{2}^{\prime}}^{2}$. Suppose that the following (i) $\sim$ (iii) holds.
  1. (i)
[…]
  2. (ii)
[…]
  3. (iii)
    $2n_{1}+3n_{2}\neq 2n_{1}^{\prime}+3n_{2}^{\prime}$.
Then we have $L_{n}\neq L_{n^{\prime}}$.
Proof.
We have
$(3n_{2})^{3}f_{n}\left(\frac{3X}{n_{2}}+\frac{n}{3}\right)=3^{6}\left(X^{3}-\dfrac{\Delta_{n}}{9}\cdot\dfrac{X}{3}-\left(\dfrac{2n_{1}}{3}+n_{2}\right)\dfrac{\Delta_{n}}{9}\cdot\dfrac{1}{27}\right),$
$(3n_{2}^{\prime})^{3}f_{n}\left(\frac{3X}{n_{2}^{\prime}}+\frac{n^{\prime}}{3}\right)=3^{6}\left(X^{3}-\dfrac{\Delta_{n^{\prime}}}{9}\cdot\dfrac{X}{3}-\left(\dfrac{2n^{\prime}_{1}}{3}+n^{\prime}_{2}\right)\dfrac{\Delta_{n^{\prime}}}{9}\cdot\dfrac{1}{27}\right).$
We can show that all prime numbers $p$ which divides $\Delta_{n}/9$ or $\Delta_{n^{\prime}}/9$ satisfy $p\equiv 0,1\pmod{3}$ (see [2, Lemma 3]). Furthermore, we have $p\neq 3$ from the assumption (ii). Therefore, both $\Delta_{n}/9$ and $\Delta_{n^{\prime}}/9$ are products of distinct prime numbers $p$ satisfying $p\equiv 1\pmod{3}$. Furthermore, we have $2n_{1}/3+n_{2}\equiv 2n_{1}^{\prime}/3+n_{2}^{\prime}\equiv 2\pmod{3}$ from (i). Since $f_{n}(X)$ and $f_{n^{\prime}}(X)$ are irreducible over $\mathbb{Q}$ by Lemma 1, both $L_{n}$ and $L_{n^{\prime}}$ are cyclic cubic fields. Using these facts and [5, Lemma 6.4.5], the roots of the right-hand sides of the two equations of (LABEL:eq:fn-lemma) give different cyclic cubic fields, and we have $L_{n}\neq L_{n^{\prime}}$. ∎
Let $L/\mathbb{Q}$ be a finite abelian extension with Galois group $G$. Leopoldt showed that the ring of integers $\mathcal{O}_{L}$ of $L$ is a free module of rank $1$ over the associated order $\mathcal{A}_{L/\mathbb{Q}}:=\{x\in\mathbb{Q}[G]\ |\ x\mathcal{O}_{L}\subset\mathcal{O}_{L}\}$ ([12, Satz 6]. [13, Theorem 2]). The following lemma is a part of a recent result of [2, Corollary 5], which is a generalization of the results [8] and [16] for the simplest cubic field.
Lemma 3.
Let $n=n_{1}/n_{2}$ be a rational number where the integers $n_{1}$ and $n_{2}$ are coprime, $\mathfrak{f}$ be the conductor of $L_{n}$. Suppose that $9||\Delta_{n}$ and $\Delta_{n}/9$ is square-free, where $\Delta_{n}=n_{1}^{2}+3n_{1}n_{2}+9n_{2}^{2}$. Put $\alpha=n_{2}\rho_{n}-n_{1}/3$. Then we have $\alpha\in{\bf e}_{\mathfrak{f}}\mathcal{O}_{L_{n}}$ for ${\bf e}_{\mathfrak{f}}=(2-\sigma-\sigma^{2})/3$ and $\alpha+1$ is a generator of $\mathcal{O}_{L_{n}}$ over $\mathcal{A}_{L_{n}/\mathbb{Q}}$, namely we have $\mathcal{O}_{L_{n}}=\mathcal{A}_{L_{n}/\mathbb{Q}}(\alpha+1)$ (see §3 for the definition of ${\bf e}_{\mathfrak{f}}\in\mathbb{Q}[G]$).
3. Structure of the units group of the associated order
Let $p$ be an odd prime number and $L/\mathbb{Q}$ a cyclic extension of degree $p$ with Galois group $G=\langle\sigma\rangle$. The conductor $\mathfrak{f}$ of $L$ should be
(3.1) $\mathfrak{f}=\begin{cases}p_{1}\cdots p_{\nu}&\text{if $p\nmid\mathfrak{f}$\, (tamely ramified)},\\ p^{2}p_{1}\cdots p_{\nu}&\text{if $p|\mathfrak{f}$\, (wildly ramified)},\end{cases}$
where $p_{1},\ldots,p_{\nu}$ are different prime numbers satisfying $p_{1}\equiv\cdots\equiv p_{\nu}\equiv 1\pmod{p}$. Let $v_{p}(x)$ denote the $p$-adic valuation of $x\in\mathbb{Q}$ for a prime number $p$. For any $m\in\mathbb{Z}_{>0}$, we put
$p(m)=\prod_{\begin{subarray}{c}p\mid m\\ p\neq 2\end{subarray}}p,\qquad q(m)=\prod_{\begin{subarray}{c}p\\ v_{p}(m)\geq 2\end{subarray}}p^{v_{p}(m)}.$
where the first product runs over all odd prime numbers $p$ dividing $m$, and the second product runs over all prime numbers $p$ that satisfy $v_{p}(m)\geq 2$. Put
$\mathcal{D}(\mathfrak{f})=\{m\in\mathbb{Z}_{>0}\,|\,p(\mathfrak{f})|m,\ m|\mathfrak{f},\ m\not\equiv 2\pmod{4}\}$.
Let $\mathscr{X}$ be the group of Dirichlet characters associated to $L$. We define a branch class of $\mathscr{X}$ for any $m\in\mathcal{D}(\mathfrak{f})$ by
$\Phi_{m}=\{\chi\in\mathscr{X}\,|\,q(\mathfrak{f}_{\chi})=q(m)\}.$
where $\mathfrak{f}_{\chi}$ is the conductor of $\chi$. We have $\mathscr{X}=\coprod_{m\in\mathcal{D}(\mathfrak{f})}\Phi_{m}$ (disjoint union). For any $\chi\in X$, let
${\bf e}_{\chi}=\frac{1}{[L:\mathbb{Q}]}\sum_{g\in G}\chi^{-1}(g)g$
be the idempotent. Furthermore, for any $m\in\mathcal{D}(\mathfrak{f})$, let
${\bf e}_{m}=\sum_{\chi\in\Phi_{m}}{\bf e}_{\chi}.$
Since the branch class $\Phi_{m}$ is closed under conjugation, we obtain ${\bf e}_{m}\in\mathbb{Q}[G]$. Since the conductor $\mathfrak{f}$ is given by (3.1) for a cyclic extension $L/\mathbb{Q}$ of degree $p\,(\neq 2)$, we have
[…]
Leopoldt ([12, Satz 6]. [13, Theorem 2]) showed that
(3.2) $\mathcal{O}_{L}=\begin{cases}\mathcal{A}_{L/\mathbb{Q}}{\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})&\text{if $p\nmid\mathfrak{f}$,}\\ \mathcal{A}_{L/\mathbb{Q}}({\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})+1)&\text{if $p|\mathfrak{f}$}.\end{cases}$
and
(3.3) $\mathcal{A}_{L/\mathbb{Q}}=\begin{cases}\mathbb{Z}[G][{\bf e}_{\mathfrak{f}}]=\mathbb{Z}[G]&\text{if $p\nmid\mathfrak{f}$,}\\ \mathbb{Z}[G][{\bf e}_{\mathfrak{f}},\,{\bf e}_{\mathfrak{f}/p}]&\text{if $p|\mathfrak{f}$}.\end{cases}$
From this result, we know that if $\eta$ is a generator of $\mathcal{A}_{L/\mathbb{Q}}$-module $\mathcal{O}_{L}$, then there exists $u\in\mathcal{A}_{L/\mathbb{Q}}^{\times}$ such that $\eta=u{\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})$ (resp. $\eta=u({\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})+1)$) if $p\nmid\mathfrak{f}$ (resp. $p|\mathfrak{f}$), and there is a one-to-one correspondence between the set of all generators of $\mathcal{O}_{L}$ and $\mathcal{A}_{L/\mathbb{Q}}^{\times}$. In this section, we consider the structure of the group $\mathcal{A}_{L/\mathbb{Q}}^{\times}$.
First, we consider the group structure of $\mathbb{Z}[G]^{\times}$. Let $U_{p}=\{u\in\mathbb{Z}[\zeta_{p}]^{\times}\,|\,u\equiv\pm 1\pmod{(1-\zeta_{p})}\}$ and $u_{k}=(1-\zeta_{p}^{k})/(1-\zeta_{p})$ for $k\in\{1,2,\cdots,(p-1)/2\}$. Let $\phi_{p}(X)=(X^{p}-1)/(X-1)\in\mathbb{Z}[X]$ be the $p$-th cyclotomic polynomial. The following lemma was proven in [1, Theorems 1.6 and 1.7] except for the injectivity.
Lemma 4 ([1]).
A group homomorphism
$\nu:\ \mathbb{Z}[G]^{\times}\longrightarrow\mathbb{Z}[\zeta_{p}]^{\times},\quad\sigma\mapsto\zeta_{p}$
is injective, $\nu(\mathbb{Z}[G]^{\times})=U_{p}$ and
$\mathbb{Z}[\zeta_{p}]^{\times}/U_{p}=\{\overline{u}_{k}\,|\,k\in\{1,2,\ldots,(p-1)/2\}\},$
where $\overline{u}_{k}=u_{k}U_{p}$.
Proof.
See [1, Theorems 1.6 and 1.7] except for the injectivity. We show that $\nu$ is injective. Put
$\psi:(\mathbb{Z}[X]/(X^{p}-1))^{\times}\longrightarrow\mathbb{Z}[\zeta_{p}]^{\times},\quad\overline{X}\mapsto\zeta_{p}.$
Since the composition of an isomorphism $\mathbb{Z}[G]^{\times}\overset{\sim}{\longrightarrow}(\mathbb{Z}[X]/(X^{p}-1))^{\times}$ and $\psi$ is $\nu$, it suffices to show that $\psi$ is injective. Let $\overline{f}\in{\rm Ker}\,\psi$. Since $f(\zeta_{p})=1,\,\phi_{p}=(X^{p}-1)/(X-1)\in\mathbb{Z}[X],\,f-1\in\mathbb{Z}[X]$ and $\phi_{p}$ is monic, $\phi_{p}$ divides f1f-1 in $\mathbb{Z}[X]$. We have $f(1)\equiv 1\pmod{p}$. Furthermore, we have $f(1)=\pm 1$ from $\overline{f}\in(\mathbb{Z}[X]/(X^{p}-1))^{\times}$ and a ring homomorphism $\mathbb{Z}[X]/(X^{p}-1)\longrightarrow\mathbb{Z},\ \overline{f}\mapsto f(1)$. From these facts and $p\neq 2$, we obtain $f(1)=1$. Therefore, X1X-1 divides f1f-1 in $\mathbb{Z}[X]$. We conclude that $X^{p}-1=(X-1)\phi_{p}$ divides f1f-1 in $\mathbb{Z}[X]$, and $\overline{f}=1$ in $(\mathbb{Z}[X]/(X^{p}-1))^{\times}$. We obtain that $\psi$ is injective. ∎
We will prove the following theorem on the structure of the group $\mathcal{A}^{\times}_{L./\mathbb{Q}}$.
Theorem 3.
Let $p$ be an odd prime number and $L/\mathbb{Q}$ a cyclic extension of degree $p$. Let $\mathfrak{f}$ denote the conductor of $L$.
  1. (1)
    If $p\nmid\mathfrak{f}$, then we have the following group isomorphism :
    $\mathcal{A}_{L/\mathbb{Q}}^{\times}\overset{\sim}{\longrightarrow}U_{p},\quad\sigma\mapsto\zeta_{p}.$
  2. (2)
    If $p|\mathfrak{f}$, then we have the following exact sequence of abelian groups :
    1{1, $\longrightarrow\{1,\,1-2{\bf e}_{\bf 1}\}\longrightarrow\mathcal{A}_{L/\mathbb{Q}}^{\times}\overset{\psi}{\longrightarrow}\mathbb{Z}[\zeta_{p}]^{\times}\longrightarrow 1$,
    where $\psi(\sigma)=\zeta_{p}$ and ${{\bf e}_{\bf 1}=\frac{1}{p}\sum_{i=0}^{p-1}\sigma^{i}}$ is the idempotent for the trivial character $\bf 1$}.
Proof.
The assertion of (1) follows from (3.3) and Lemma 4. We prove (2). First, we will prove that the map $\psi$ is surjective. Let $\alpha\in\mathbb{Z}[\zeta_{p}]^{\times}$. By Lemma 4, there exists $u\in U_{p}$ and $k\in\{1,2,\ldots,(p-1)/2\}$ such that $\alpha=u_{k}u$. Let $\beta\in\mathbb{Z}[\zeta_{p}]^{\times}$ satisfy $\alpha\beta=1$ and write $\beta=u_{\ell}v,\,v\in U_{p}$ and $\ell\in\{1,2,\ldots,(p-1)/2\}$. We can write $\alpha=f(\zeta_{p})$ and $\beta=g(\zeta_{p})$ for $f,g\in\mathbb{Z}[X]$. Since $f(\zeta_{p})g(\zeta_{p})=\alpha\beta=1$, we have
(3.4) $fg\equiv 1\pmod{(\phi_{p})}\quad{\rm in}\ \mathbb{Z}[X].$
On the other hand, since $\alpha=u_{k}u\equiv\pm u_{k}\equiv\pm k\pmod{(1-\zeta_{p})}$ and $\beta=u_{\ell}v\equiv\pm u_{\ell}\equiv\pm\ell\pmod{(1-\zeta_{p})}$, we obtain $f\equiv a\pmod{(\phi_{p},X-1)},\ a:=\pm k$ and $g\equiv b\pmod{(\phi_{p},X-1)},\ b:=\pm\ell$. Let $f_{1},g_{1}\in\mathbb{Z}[X]$ satisfy
(3.5) $f\equiv a+f_{1}\phi_{p}\qquad\pmod{(X-1)}$,
$g\equiv\,b+g_{1}\phi_{p}\qquad\pmod{(X-1)}$.
From (3.4) and (3.5), we have $1\equiv fg\equiv ab\pmod{(\phi_{p},X-1)}$, and hence $ab\equiv 1\pmod{p}$. Let $c\in\mathbb{Z}$ satisfy $ab=1+pc$. Define $f_{2},g_{2}\in\mathbb{Q}[X]$ by
$f_{2}=f-f_{1}\phi_{p}-\frac{a+1}{p}\,\phi_{p},$
$g_{2}=g-g_{1}\phi_{p}-\frac{b+1}{p}\,\phi_{p}.$
By (3.4), we have
$f_{2}g_{2}\equiv fg\equiv 1\pmod{(\phi_{p})}\quad{\rm in}\ \mathbb{Q}[X],$
and hence
(3.6) $\phi_{p}\,|\,(f_{2}g_{2}-1)\qquad{\rm in}\ \mathbb{Q}[X].$
On the other hand, we have by (3.5)
$f_{2}g_{2}\equiv 1\pmod{(X-1)}\qquad{\rm in}\ \mathbb{Q}[X],$
and hence
(3.7) $(X-1)\,|\,(f_{2}g_{2}-1)\quad{\rm in}\ \mathbb{Q}[X].$
From (3.6) and (3.7), we conclude that $X^{p}-1=(X-1)\phi_{p}$ divides $f_{2}g_{2}-1$ in $\mathbb{Q}[X]$, and $\overline{f}_{2}\overline{g}_{2}=1$ in $\mathbb{Q}[X]/(X^{p}-1)$. Define $y,z\in\mathcal{A}_{L/\mathbb{Q}}$ by
$y=f(\sigma)-f_{1}(\sigma)p{\bf e}_{\bf 1}-(a+1){\bf e}_{\bf 1},$
$z=g(\sigma)-g_{1}(\sigma)p{\bf e}_{\bf 1}-(b+1){\bf e}_{\bf 1}.$
The image of $yz$ by the map $\mathcal{A}_{L/\mathbb{Q}}\hookrightarrow\mathbb{Q}[G]\simeq\mathbb{Q}[X]/(X^{p}-1),\ \sigma\mapsto\overline{X}$ is $\overline{f}_{2}\overline{g}_{2}=1$, we obtain $yz=1$, and hence $y,z\in\mathcal{A}_{L/\mathbb{Q}}^{\times}$. Since $\psi({\bf e}_{\bf 1})=0$, we have $\alpha=f(\zeta_{p})=\psi(y),\ y\in\mathcal{A}_{L/\mathbb{Q}}^{\times}$, and hence $\alpha\in\psi(\mathcal{A}_{L/\mathbb{Q}}^{\times})$. We obtain that $\psi$ is surjective.
Next, we show that ${\rm Ker}\,\psi=\{1,1-2{\bf e}_{\bf 1}\}$. From (3.3) and ${\bf e}_{\mathfrak{f}}+{\bf e}_{\mathfrak{f}/p}=1,\ {\bf e}_{\mathfrak{f}/p}={\bf e}_{\bf 1}$, we have
$\mathcal{A}_{L/\mathbb{Q}}=\mathbb{Z}[G][{\bf e}_{\mathfrak{f}},{\bf e}_{\mathfrak{f}/p}]=\mathbb{Z}[G][{\bf e}_{\bf 1}].$
Furthermore, since for any $x\in\mathbb{Z}[G]$ there exists $a\in\mathbb{Z}$ such that $x{\bf e}_{\bf 1}=a{\bf e}_{\bf 1}$ and ${\bf e}_{\bf 1}^{2}={\bf e}_{\bf 1}$, we obtain $\mathcal{A}_{L/\mathbb{Q}}=\{x+a{\bf e}_{\bf 1}\,|\,x\in\mathbb{Z}[G],\ a\in\mathbb{Z}\}$. First, we show ${\rm Ker}\,\psi=\{1+a{\bf e}_{\bf 1}\,|\,a\in\mathbb{Z}\}\cap\mathcal{A}_{L/\mathbb{Q}}^{\times}$. Let $\widetilde{\psi}:\mathcal{A}_{L/\mathbb{Q}}^{\times}\overset{\psi}{\longrightarrow}\mathbb{Z}[\zeta_{p}]^{\times}\overset{\sim}{\longrightarrow}(\mathbb{Z}[X]/(\phi_{p}))^{\times}$. It is enough to show ${\rm Ker}\,\widetilde{\psi}=\{1+a{\bf e}_{\bf 1}\,|\,a\in\mathbb{Z}\}\cap\mathcal{A}_{L/\mathbb{Q}}^{\times}$. We will prove ${\rm Ker}\,\widetilde{\psi}\subseteq\{1+a{\bf e}_{\bf 1}\,|\,a\in\mathbb{Z}\}\cap\mathcal{A}_{L/\mathbb{Q}}^{\times}$ since the opposite inclusion is trivial. Let $\alpha=x+a{\bf e}_{\bf 1}\in{\rm Ker}\,\widetilde{\psi},\,x\in\mathbb{Z}[G],\,a\in\mathbb{Z}$. we can write $x=f(\sigma)$ for $f(X)\in\mathbb{Z}[X]$. Since
$1=\widetilde{\psi}(\alpha)=\widetilde{\psi}(f(\sigma)+a{\bf e}_{\bf 1})=\overline{f(X)},$
we have $f(X)\equiv 1\pmod{(\phi_{p})}$ in $\mathbb{Z}[X]$. Let $g\in\mathbb{Z}[X]$ satisfy $f=1+g\Phi_{p}$, and $c\in\mathbb{Z}$ satisfy $g(\sigma){\bf e}_{\bf 1}=c{\bf e}_{\bf 1}$. Then we have
$\alpha=x+a{\bf e}_{\bf 1}=f(\sigma)+a{\bf e}_{\bf 1}=1+g(\sigma)p{\bf e}_{\bf 1}+a{\bf e}_{\bf 1}=1+(cp+a){\bf e}_{\bf 1},$
and hence $\alpha\in\{1+a{\bf e}_{\bf 1}\,|\,a\in\mathbb{Z}\}\cap\mathcal{A}_{L/\mathbb{Q}}^{\times}$. We obtain ${\rm Ker}\,\psi=\{1+a{\bf e}_{\bf 1}\,|\,a\in\mathbb{Z}\}\cap\mathcal{A}_{L/\mathbb{Q}}^{\times}$, and to show ${\rm Ker}\,\psi=\{1,1-2{\bf e}_{\bf 1}\}$, we show $\{1+a{\bf e}_{\bf 1}\,|\,a\in\mathbb{Z}\}\cap\mathcal{A}_{L/\mathbb{Q}}^{\times}=\{1,1-2{\bf e}_{\bf 1}\}$. We have $1-2{\bf e}_{\bf 1}\in\{1+a{\bf e}_{\bf 1}\,|\,a\in\mathbb{Z}\}\cap\mathcal{A}_{L/\mathbb{Q}}^{\times}$ since $(1-2{\bf e}_{\bf 1})^{2}=1$. Conversely, let $1+a{\bf e}_{\bf 1}\in\mathcal{A}_{L/\mathbb{Q}}^{\times},\ a\in\mathbb{Z}$. We will show that $a\in\{0,-2\}$. Since $\{1+a{\bf e}_{\bf 1}\,|\,a\in\mathbb{Z}\}\cap\mathcal{A}_{L/\mathbb{Q}}^{\times}={\rm Ker}\,\psi$ is a subgroup of $\mathcal{A}_{L/\mathbb{Q}}^{\times}$, there exists $1+b{\bf e}_{\bf 1}\in\mathcal{A}_{L/\mathbb{Q}}^{\times},\ b\in\mathbb{Z}$ satisfying
$1=(1+a{\bf e}_{\bf 1})(1+b{\bf e}_{\bf 1})=1+(ab+a+b){\bf e}_{\bf 1}.$
From this equality, we obtain $ab+a+b=0$. The pair $(a,b)=(0,0)$ satisfies $ab+a+b=0$. We assume that $(a,b)\neq(0,0)$. Since a(1+b)=ba(1+b)=-b, we have $b|a$. Let $t\in\mathbb{Z}$ satisfy $a=bt$. Then we have t(1+b)=1t(1+b)=-1. Since $t,b\in\mathbb{Z}$ and $b\neq 0$, we conclude that b=2b=-2 and $t=1$, and hence (a,b)=(2,2)(a,b)=(-2,-2). We obtain ${\rm Ker}\,\psi=\{1,\,1-2{\bf e}_{\bf 1}\}$. ∎
Corollary 1.
If $p=3$, then we have
$\mathcal{A}_{L/\mathbb{Q}}^{\times}=\begin{cases}\langle-1\rangle\times\langle\sigma\rangle=\{\pm 1,\pm\sigma,\pm\sigma^{2}\}&{\rm if}\ 3\nmid\mathfrak{f},\\ \langle 1-2{\bf e}_{\bf 1}\rangle\times\langle-1\rangle\times\langle\sigma\rangle=\{\pm 1,\pm\sigma,\pm\sigma^{2},\pm(1-2{\bf e}_{\bf 1}),\pm\sigma(1-2{\bf e}_{\bf 1}),\pm\sigma^{2}(1-2{\bf e}_{\bf 1})\}&{\rm if}\ 3|\mathfrak{f}.\end{cases}$
Proof.
The assertion follows from Theorem 3 and $\mathbb{Z}[\zeta_{3}]^{\times}=U_{3}=\{\pm 1,\pm\zeta_{3},\pm\zeta_{3}^{2}\}$. ∎
4. Proof of the theorem
In this section, we give the proof of Theorem 2. First, we have $n_{1},n_{2}\in\mathbb{Z}$ since $M\equiv N\pmod{2}$ from $4\mathfrak{f}=M^{2}+27N^{2}$. We show that $n_{1}$ and $n_{2}$ are coprime. Let $p$ be a prime number which divides both $n_{1}$ and $n_{2}$. Since $4\mathfrak{f}=M^{2}+27N^{2}$ and $2\nmid\mathfrak{f}$, either $n_{1}$ or $n_{2}$ is not divisible by $2$ (note that $M\equiv N\equiv 1\pmod{2}$ or $M\not\equiv N\pmod{4}$ holds), and hence $p\neq 2$. Furthermore, we have $p\neq 3$ since $n_{2}=N\not\equiv 0\pmod{3}$. Since $p$ divides both $n_{1}=(M-3N)/2$ and $n_{2}=N$, it follows that $p$ divides $M$. This is a contradiction since $4\mathfrak{f}=M^{2}+27N^{2}$, $\mathfrak{f}/9$ is square-free and $p\neq 2,3$. Therefore, $n_{1}$ and $n_{2}$ are coprime. Since $4\mathfrak{f}=M^{2}+27N^{2}=4(n_{1}^{2}+3n_{1}n_{2}+9n_{2}^{2})$, we have $\mathfrak{f}=n_{1}^{2}+3n_{2}n_{2}+9n_{2}^{2}$, and hence $3|n_{1}$. It follows that $f_{n}(X)$ for $n=n_{1}/n_{2}$ is irreducible over $\mathbb{Q}$ by Lemma 1. Let $t=n_{1}/3\in\mathbb{Z}$. From $n_{1}=(M-3N)/2$ and $M=3M_{0}$, we have $2t+n_{2}=M_{0}\equiv 2\pmod{3}$. and hence $t\not\equiv n_{2}\pmod{3}$. It follows that the conductor of $L_{n}$ is $\mathfrak{f}$ ([2, Corollary 1] and $D_{L_{n}}=\mathfrak{f}^{2}$ where $D_{L_{n}}$ is the discriminant of $L_{n}$). We have already proved (1), (2), (3) of the theorem.
Next, we prove the remaining (4) and (5). From Lemma 3, $\alpha+1$ is a generator of $\mathcal{O}_{L_{n}}$ over $\mathcal{A}_{L_{n}/\mathbb{Q}}$ for $\alpha:=n_{2}\rho_{n}-n_{1}/3\in{\bf e}_{\mathfrak{f}}\mathcal{O}_{L_{n}}$. Let $\alpha^{\prime}=\sigma(\alpha)$ and $\alpha^{\prime\prime}=\sigma^{2}(\alpha)$. Since
$\mathcal{A}_{L_{n}/\mathbb{Q}}^{\times}=\{\pm 1,\pm\sigma,\pm\sigma^{2},\pm(1-2{\bf e}_{\bf 1}),\pm\sigma(1-2{\bf e}_{\bf 1}),\pm\sigma^{2}(1-2{\bf e}_{\bf 1})\}$
from Corollary 1, there are 12 generators of $\mathcal{O}_{L_{n}}$ over $\mathcal{A}_{L_{n}/\mathbb{Q}}$ which are given by $u(\alpha+1)$ for $u\in\mathcal{A}^{\times}_{L_{n}/\mathbb{Q}}$. Since ${\bf e}_{\mathfrak{f}}\,{\bf e}_{\bf 1}=0$ and $\alpha\in{\bf e}_{\mathfrak{f}}\,\mathcal{O}_{L_{n}}$, we have $(1-2{\bf e}_{\bf 1})(\alpha+1)=\alpha-1$. Therefore, the 12 generators of $\mathcal{O}_{L_{n}}$ are
(4.1) $\pm\alpha\pm 1,\quad\pm\alpha^{\prime}\pm 1,\quad\pm\alpha^{\prime\prime}\pm 1\quad\text{(any double sign)}.$
On the other hand, from (3.2), we know that $\eta_{0}+1$ is a generator where $\eta_{0}={\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L_{n})}(\zeta_{\mathfrak{f}})$, and hence $\eta_{0}+1$ is equal to one of the 12 generators of (4.1). Since ${\bf e}_{\mathfrak{f}}\,\eta_{0}=\eta_{0},\ {\bf e}_{\mathfrak{f}}\,1=0$, $\{\eta_{0},\eta_{1},\eta_{2}\}$ must be $\{\alpha,\alpha^{\prime},\alpha^{\prime\prime}\}$ or $\{-\alpha,-\alpha^{\prime},-\alpha^{\prime\prime}\}$. Let $h(X)=(X-\alpha)(X-\alpha^{\prime})(X-\alpha^{\prime\prime})$ be the minimal polynomial of $\alpha$. Since $\rho_{n},\rho_{n}^{\prime}$ and $\rho_{n}^{\prime\prime}$ are roots of $f_{n}(X)$ and $\alpha=n_{2}\rho_{n}-n_{1}/3$, we have
(4.2) $n_{2}^{3}f_{n}(X)=h\left(n_{2}X-\frac{n_{1}}{3}\right).$
Let $P(X)=(X-\eta_{0})(X-\eta_{1})(X-\eta_{2})$ be the period polynomial. Since $\{\eta_{0},\eta_{1},\eta_{2}\}=\{\alpha,\alpha^{\prime},\alpha^{\prime\prime}\}$ or $\{-\alpha,-\alpha^{\prime},-\alpha^{\prime\prime}\}$ and the coefficients of $x^{2}$ are zero, the difference of two polynomials $h(X)$ and $P(X)$ is only the sign of the constant term. To determine the sign, we calculate the values of $\eta_{0}\eta_{1}\eta_{2}$ and $\alpha\alpha^{\prime}\alpha^{\prime\prime}$ modulo $3$. Let $\mathscr{X}=\langle\chi\rangle$ be the group of Dirichlet characters associated to $L_{n}$ where $\chi=\chi_{3^{2}}\chi_{p_{1}}\cdots\chi_{p_{\nu}}$ and $\chi_{m}$ is the Dirichlet character of conductor $m$. We define the Gaussian sum $\tau(\chi)$ for the character $\chi$ of conductor $\mathfrak{f}$ is
$\tau(\chi)=\sum_{a\in(\mathbb{Z}/\mathfrak{f}\mathbb{Z})^{\times}}\chi(a)\zeta_{\mathfrak{f}}^{a}.$
We have $\tau(\chi)=\eta_{0}+\zeta_{3}\eta_{1}+\zeta_{3}^{2}\eta_{2}$ or $\eta_{0}+\zeta_{3}^{2}\eta_{1}+\zeta_{3}\eta_{2}$ for the primitive third root of unity $\zeta_{3}=e^{2\pi i/3}$. Since $\eta_{0}+\eta_{1}+\eta_{2}={\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L_{n}}(\zeta_{\mathfrak{f}})=\mu(\mathfrak{f})=0$, we have $\tau(\chi)+\overline{\tau(\chi)}=3\eta_{0}$, and hence we obtain
$(3\eta_{0})^{3}=\tau(\chi)^{3}+\overline{\tau(\chi)}^{3}+9\eta_{0}\tau(\chi)\,\overline{\tau(\chi)}.$
Furthermore, since $\tau(\chi)\,\overline{\tau(\chi)}=\mathfrak{f}=3^{2}p_{1}\cdots p_{\nu}$, it conclude that
(4.3) $(3\eta_{0})^{3}=\tau(\chi)^{3}+\overline{\tau(\chi)}^{3}+3^{4}p_{1}\cdots p_{\nu}\eta_{0}.$
Since $\chi=\chi_{3^{2}}\chi_{p_{1}}\cdots\chi_{p_{\nu}}$, we have
(4.4) $\tau(\chi)^{3}=\tau(\chi_{3^{2}})^{3}\tau(\chi_{p_{1}})^{3}\cdots\tau(\chi_{p_{\nu}})^{3}.$
By direct calculation, we have
(4.5) $\tau(\chi_{3^{2}})=\begin{cases}27\zeta_{3}&\text{if $\chi_{3^{2}}(a)=\zeta_{3}^{(a^{2}-1)/3}$,}\\ 27\zeta_{3}^{-1}&\text{if $\chi_{3^{2}}(a)=\zeta_{3}^{-(a^{2}-1)/3}$,}\end{cases}$
(4.6) $\tau(\chi_{p_{i}})=\sum_{a\in(\mathbb{Z}/p_{i}\mathbb{Z})^{\times}}\chi_{p_{i}}(a)\zeta_{p_{i}}^{a}\equiv-1\pmod{(1-\zeta_{3})}\qquad\text{in}\ \mathbb{Z}[\zeta_{\mathfrak{f}}]$
for $i\in\{1,\ldots,\nu\}$. From (4.3), (4.4), (4.5) and (4.6), we obtain
$(3\eta_{0})^{3}$ $\equiv 2\times(-1)^{\nu}\times 27+3^{4}p_{1}\cdots p_{\nu}\eta_{0}$
$\equiv 2\times(-1)^{\nu}\times 27\pmod{27(1-\zeta_{3})}\qquad\text{in}\ \mathbb{Z}[\zeta_{\mathfrak{f}}].$
We conclude that $\eta_{0}^{3}\equiv 2\times(-1)^{\nu}\equiv(-1)^{\nu+1}\pmod{(1-\zeta_{3})}$ and hence $(\eta_{0}\eta_{1}\eta_{2})^{3}\equiv(-1)^{\nu+1}\pmod{(1-\zeta_{3})}$. Since $\eta_{0}\eta_{1}\eta_{2}\in\mathbb{Z}$, we have $(\eta_{0}\eta_{1}\eta_{3})^{3}\equiv(-1)^{\nu+1}\pmod{3}$ and hence
(4.7) $\eta_{0}\eta_{1}\eta_{2}\equiv(-1)^{\nu+1}\pmod{3}$.
On the other hand, we obtain
(4.8) $\alpha\alpha^{\prime}\alpha^{\prime\prime}$ $=\left(n_{2}\rho_{n}-\frac{n_{1}}{3}\right)\left(n_{2}\rho_{n}^{\prime}-\frac{n_{1}}{3}\right)\left(n_{2}\rho_{n}^{\prime\prime}-\frac{n_{1}}{3}\right)$
$=\frac{1}{27}(n_{1}^{2}+3n_{1}n_{2}+9n_{2}^{2})(2n_{1}+3n_{2})$
$=\frac{1}{27}\mathfrak{f}M=p_{1}\cdots p_{\nu}M_{0}\equiv-1\pmod{3}$.
From (4.7) and (4.8), we have
$\{\eta_{0},\eta_{1},\eta_{2}\}=\{(-1)^{\nu}\alpha,(-1)^{\nu}\alpha^{\prime},(-1)^{\nu}\alpha^{\prime\prime}\}=\{\mu(\mathfrak{f}/9)\alpha,\mu(\mathfrak{f}/9)\alpha^{\prime},\mu(\mathfrak{f}/9)\alpha^{\prime\prime}\}.$
Therefore, we have
(4.9) $h(X)=\mu(\mathfrak{f}/9)P(\mu(\mathfrak{f}/9)X).$
From (4.2) and (4.9), we obtain (4) of the theorem, and (5) follows from (4).
Finally, let $(M^{\prime},N^{\prime})$ be another pair satisfying (1.5) and put $n_{1}^{\prime}=(M^{\prime}-3N^{\prime})/2$ and $n_{2}^{\prime}=N^{\prime}$. Since $2n_{1}+3n_{2}=M\neq M^{\prime}=2n_{1}^{\prime}+3n_{2}^{\prime}$, we have $L_{n}\neq L_{n^{\prime}}$ by Lemma 2. Since there are exactly $2^{\nu-1}$ pairs $(M,N)$ which satisfy (1.5) ([5, p.3̇42, 343]) and there are exactly $2^{\nu-1}$ cubic fields with conductor $\mathfrak{f}$, any cyclic cubic field with conductor $\mathfrak{f}$ must coincide with $L_{n}$ for $n=n_{1}/n_{2}$ where $n_{1}$ and $n_{2}$ are defined by such a pair $(M,N)$.
5. Examples
We consider cyclic cubic fields with conductor $\mathfrak{f}=9\times 7\times 13$. All the pairs $(M,N)$ satisfying (1.5) are $(M,N)=(-3\cdot 19,1),(3\cdot 17,5),(3\cdot 8,10),(-3,11)$, and the corresponding pair $(n_{1},n_{2})$ are (30,1),(18,5),(3,10),(18,11)(-30,1),(18,5),(-3,10),(-18,11) in order. Table 1 shows the Gaussian periods of $L_{n}$, the period polynomials and Shanks’ cubic polynomials for each $n=n_{1}/n_{2}$.
Table 1. $\mathfrak{f}=9\times 7\times 13$
$(n_{1},n_{2})$ $\{\eta_{0},\eta_{1},\eta_{2}\}$ $P(X)$    $f_{n}(X)$
(30,1)(-30,1) $\{\rho+10,\,\rho^{\prime}+10,\,\rho^{\prime\prime}+10\}$ $X^{3}+30X^{2}+27X-1$ $X^{3}-273X+1729$
[…]

### v3

@@
 Leopoldt ([12, Satz 6]. [13, Theorem 2]) showed that
-(3.2) $\mathcal{O}_{L}=\begin{cases}\mathcal{A}_{L/\mathbb{Q}}{\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})&\text{if $p\nmid\mathfrak{f}$,}\\ \mathcal{A}_{L/\mathbb{Q}}({\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})+1)&\text{if $p|\mathfrak{f}$}.\end{cases}$
+(3.2) $\mathcal{O}_{L}=\begin{cases}\mathcal{A}_{L/\mathbb{Q}}{\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L}(\zeta_{\mathfrak{f}})&\text{if $p\nmid\mathfrak{f}$,}\\ \mathcal{A}_{L/\mathbb{Q}}({\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L}(\zeta_{\mathfrak{f}})+1)&\text{if $p|\mathfrak{f}$}.\end{cases}$
 and
 (3.3) $\mathcal{A}_{L/\mathbb{Q}}=\begin{cases}\mathbb{Z}[G][{\bf e}_{\mathfrak{f}}]=\mathbb{Z}[G]&\text{if $p\nmid\mathfrak{f}$,}\\ \mathbb{Z}[G][{\bf e}_{\mathfrak{f}},\,{\bf e}_{\mathfrak{f}/p}]&\text{if $p|\mathfrak{f}$}.\end{cases}$
-From this result, we know that if $\eta$ is a generator of $\mathcal{A}_{L/\mathbb{Q}}$-module $\mathcal{O}_{L}$, then there exists $u\in\mathcal{A}_{L/\mathbb{Q}}^{\times}$ such that $\eta=u{\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})$ (resp.
-$\eta=u({\rm Tr}_{\mathbb{Q}(\mathfrak{f})/L}(\zeta_{\mathfrak{f}})+1)$) if $p\nmid\mathfrak{f}$ (resp.
+From this result, we know that if $\eta$ is a generator of $\mathcal{A}_{L/\mathbb{Q}}$-module $\mathcal{O}_{L}$, then there exists $u\in\mathcal{A}_{L/\mathbb{Q}}^{\times}$ such that $\eta=u{\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L}(\zeta_{\mathfrak{f}})$ (resp.
+$\eta=u({\rm Tr}_{\mathbb{Q}(\zeta_{\mathfrak{f}})/L}(\zeta_{\mathfrak{f}})+1)$) if $p\nmid\mathfrak{f}$ (resp.
 $p|\mathfrak{f}$), and there is a one-to-one correspondence between the set of all generators of $\mathcal{O}_{L}$ and $\mathcal{A}_{L/\mathbb{Q}}^{\times}$.
@@
 All the pairs $(M,N)$ satisfying (1.5) are $(M,N)=(-3\cdot 19,1),(3\cdot 17,5),(3\cdot 8,10),(-3,11)$, and the corresponding pair $(n_{1},n_{2})$ are (30,1),(18,5),(3,10),(18,11)(-30,1),(18,5),(-3,10),(-18,11) in order.
-Table 1 shows the Gaussian periods of $L_{n}$, the period polynomials and Shanks’ cubic polynomials for each $n=n_{1}/n_{2}$.
+Table 1 shows the Gaussian periods of $L_{n}$, Shanks’ cubic polynomials, and the period polynomials for each $n=n_{1}/n_{2}$.
+Put $\rho=\rho_{n}$.
 Table 1. $\mathfrak{f}=9\times 7\times 13$
-$(n_{1},n_{2})$ $\{\eta_{0},\eta_{1},\eta_{2}\}$ $P(X)$    $f_{n}(X)$
+$(n_{1},n_{2})$ $\{\eta_{0},\eta_{1},\eta_{2}\}$ $f_{n}(X)$    $P(X)$
 (30,1)(-30,1) $\{\rho+10,\,\rho^{\prime}+10,\,\rho^{\prime\prime}+10\}$ $X^{3}+30X^{2}+27X-1$ $X^{3}-273X+1729$